# arrows,decorations.markings,shapes,matrix,positioning, calc, fit,decorations.pathreplacing,angles,quotes,backgrounds

### v1

@@ Appendix C Human Study Interface
Refer to caption Figure 8. Example task question without explanations. Refer to caption Figure 9. Example attention check question. Refer to caption Figure 10. Bias disclosure showing the demographic parity of the model in phase 1. Refer to caption Figure 11. Correlation disclosure showing the relationship between university and gender in our synthetic data. Refer to caption Figure 12. In proxy conditions where participants are not given correlation disclosure. They are instead given this screen explaining that proxies can general, without mentioning the relationship in our data. Refer to caption Figure 13. Comprehension check screen testing both understanding of bias disclosure (Figure 10) and correlation disclosure (Figure 11). The correlation disclosure question is only shown in proxy conditions where the participants are given correlation disclosure. Refer to caption Figure 14. Initial trust survey given before the task is introduced. Refer to caption Figure 15. Example post-task survey. This is the version that is shown after phase 2 of proxy conditions. In protected conditions and after phase 1 of proxy conditions, the question about which feature might have lead to gender bias is omitted. Refer to caption Figure 16. Participant demographic questions.
(a) Fairness Perceptions (§5.1.2) (b) Decision-Making Fairness (§5.1.1) (c) Decision-Making Quality (§5.1.3) Figure 7. Effect of disclosures with explanations on various metrics when bias stems from usage of a protected vs proxy feature. The marks show the average and standard error of the given metric across participants in the given condition.
(a) Fairness Perceptions (§5.1.2) (b) Decision-Making Fairness (§5.1.1) (c) Decision-Making Quality (§5.1.3) Figure 6. Effect of disclosures without explanations on various metrics when bias stems from usage of a protected vs proxy feature. The marks show the average and standard error of the given metric across participants in the given condition.
(a) Fairness Perceptions (§5.1.2) (b) Decision-Making Fairness (§5.1.1) (c) Decision-Making Quality (§5.1.3) Figure 5. Effect of explanations alone on various metrics when bias stems from usage of a protected vs proxy feature. The marks show the average and standard error of the given metric across participants in the given condition.
Figure 2. Order of study phases.

### v2

@@ Appendix C Human Study Interface
Figure 8. Example task question without explanations.
Figure 9. Example attention check question.
Figure 10. Bias disclosure showing the demographic parity of the model in phase 1.
Figure 11. Correlation disclosure showing the relationship between university and gender in our synthetic data.
Figure 12. In proxy conditions where participants are not given correlation disclosure. They are instead given this screen explaining that proxies can general, without mentioning the relationship in our data.
Figure 13. Comprehension check screen testing both understanding of bias disclosure (Figure 10) and correlation disclosure (Figure 11). The correlation disclosure question is only shown in proxy conditions where the participants are given correlation disclosure.
Figure 14. Initial trust survey given before the task is introduced.
Figure 15. Example post-task survey. This is the version that is shown after phase 2 of proxy conditions. In protected conditions and after phase 1 of proxy conditions, the question about which feature might have lead to gender bias is omitted.
Figure 16. Participant demographic questions.